# IEEE 802.11be Wi-Fi 7: Feature Summary and Performance Evaluation

### v2

@@ Abstract
While the pace of commercial scale application of Wi-Fi 6 accelerates, the IEEE 802.11 Working Group is about to complete the development of a new amendment standard IEEE 802.11be – Extremely High Throughput (EHT), also known as Wi-Fi 7, which can be used to meet the demand for the throughput of 4K/8K videos up to tens of Gbps and low-latency video applications such as virtual reality (VR) and augmented reality (AR). Wi-Fi 7 not only scales Wi-Fi 6 with doubled bandwidth, but also supports real-time applications, which brings revolutionary changes to Wi-Fi. In this article, we start by introducing the main objectives and timeline of Wi-Fi 7 and then list the latest key techniques which promote the performance improvement of Wi-Fi 7. Finally, we validate the most critical objectives of Wi-Fi 7 – the potential up to 30 Gbps throughput and lower latency. System-level simulation results suggest that by combining the new techniques, Wi-Fi 7 achieves 30 Gbps throughput and lower latency than Wi-Fi 6.
@@ I Introduction
In this article, we introduce the application scenarios of the next-generation Wi-Fi and the main objectives of IEEE 802.11be, and summarize the recently launched standardization timeline of Wi-Fi 7. We delved into the latest concrete features to be adopted in the 802.11be amendment and detailed the implementations and benefits of each feature. In addition, we discuss two important and appealing issues – latency and throughput. In particular, we provide system-level simulation results to show that the features defined by 802.11be can achieve the throughput of 30 Gbps and improve latency performance. Altogether, this article aims to summarize Wi-Fi 7 and provide a guide for researchers and general audiences.
@@ II Wi-Fi 7 Progress
The standardization timeline of Wi-Fi 7 is illustrated in Fig. 1. The IEEE 802.11 WG approved the formation of a Topic Interest Group (TIG) in May 2018, and further established a Study Group (SG) and a Task Group (TG) in July 2018 and May 2019 respectively. The primary objective of these groups is to ensure the rapid and stable progress of 802.11be EHT.
In September 2020, 802.11be draft 0.1 was released, followed by draft 1.0 and draft 2.0 in May 2021 and May 2022, respectively. Just a few months ago, 802.11be draft 3.0 was released in January 2023, and 802.11be draft 4.0 was released in July 2023, which is the latest version available to the public. According to schedule, final amendment will be released in December 2024, and Wi-Fi 7 will make great progress in practical application. The 802.11be amendment spans five years and is expected to provide significant performance enhancements for Wi-Fi 7 users in all respects.
@@
Figure 1: Illustration of Wi-Fi 7 standardization timeline.
@@ III Wi-Fi 7 Features
Spectrum is the cornerstone of wireless communication, and the emergence of new generation of wireless communication technologies usually accompanies the use of new spectrum bands. However, the obtainable bandwidth over 2.4 GHz and 5 GHz frequency bands cannot meet the throughput and latency requirements of emerging service such as AR/VR and online gaming [2]. Therefore, the 6 GHz band is used to improve the Wi-Fi peak throughput. Researchers have discussed the most efficient way to utilize 1.2 GHz accessible unlicensed spectrum and introduced 320 MHz channels, which doubles the maximum nominal throughput compared to 802.11ax.
@@ III-E Multi-Link Operations (MLO)
With the advancement of technology, the available Wi-Fi channels are becoming wider and wider. However, for some reasons, the usage of wide channels is not efficient. For example, under the previous protocols, if the primary 20 MHz subchannel controlling the channel access is busy, the entire wide channel is blocked and the remaining subchannels are idle. In fact, a wide channel consists of multiple sub-bands with specific frequencies, so the properties of various parts of a wide channel are different, which means that they may require different parameters of channel access.
The MLO is an important new feature of 802.11be. Through MLO, APs and STAs will be provided with the capabilities to transmit and receive data from the same traffic flow over multiple radio interfaces. Its adoption represents the development tendency towards multi-link communications. TGbe has made changes in architecture, management functions, and power save to implement MLO and utilizes all available spectrum resources in 2.4 GHz, 5 GHz and 6 GHz. Next, we explore the key features of MLO.
@@ III-E1 Architecture
802.11be defines traditional APs and STAs as multi-link devices (MLDs). Both AP MLDs and STA MLDs are single devices with multiple physical radio interfaces. However, it is noteworthy that there is only one MAC address per MLD, and individual parameters for each interface are not lost. Fig. 2 shows the MLD architecture, multiple connections can be set up between APs and STAs. MLD can be considered as a collaborative work scheme. This scheme allows retransmission of packets on any link, which simplifies dynamic link switching and packet reassembly.
MLD is divided into two different transmission modes: simultaneous TX and RX (STR) mode and non-STR (NSTR) mode [11]. STR MLD can transmit frames asynchronously over multiple links. This mode can be applied to both AP MLD and non-AP MLD and keep independent channel access in different links. STR MLD can achieve higher throughput, but power consumption may increase due to simultaneous operation of multiple interfaces. Therefore, this mode must be followed by a power save mechanism, which will be described later. NSTR MLD does not allow information to be transmitted over an idle interface while they are receiving through another and can be applied to non-AP MLD. This mode can avoid in-device coexistence (IDC) interference problems, but at the cost of reducing throughput.
@@ III-E3 Power Save
Since the most major source of traffic in the information age is portable mobile devices, issues related to power consumption must be carefully considered. The use of MLD results in the presence of at least two devices at the same time, which consumes twice more energy. When the traffic is light, the device does not need to continuously listen to more than one link. The basic power management mechanism defines two working modes for Wi-Fi devices: active mode and power save (PS) mode [12]. In active mode, the device normally transmits and receives frames. In PS mode, the device is in a dozing state.
The most superior power saving mechanism designed by 802.11be is target wake time (TWT). TWT originated from 802.11ah and was further improved in 802.11ax. The key part of TWT is the initial negotiation mechanism, in which the TWT session parameters are determined [10]. Through the TWT, the STA may exchange frames with the AP during a period of waking up (TWT service periods), and keep dozing out of the TWT service periods. Under the MLO architecture, an MLD may negotiate individual TWT agreements with a peer MLD. Note that the individual TWT agreement is negotiated between the STAs affiliated with the MLDs that are operating on an enabled link and is not negotiated between two MLDs.
@@ III-F Low-Latency Operations
TABLE II: Simulation settings with three or four R1 feature
@@
Figure 3: Simulation results of 802.11be throughput (with three or four R1 features).
@@ IV-A PAR throughput verification
In the validation process for achieving 30 Gbps throughput, four release 1 (R1) features were considered to meet the objectives outlined in the PAR for the project. These features included multi-link support (with 2 links and STR transmission mode), 4K-QAM modulation, aggregation of 1024 MPDU aggregation per A-MPDU, and a 320 MHz bandwidth. These features collectively ensure a guaranteed 30 Gbps throughput for both single user (SU) and multi-user (MU) scenarios. In the simulation scenario, each BSS consists of an AP MLD and $N$ non-AP MLDs placed around it and the throughput is measured at MAC data service access point (SAP). In terms of simulation setup parameters, the number of non-AP MLD is 8 and the size of area is 20 m $\times$ 20 m. The TX power of AP MLD is 20 dBm and the TX power of non-AP MLD is 15 dBm. The number of links is 2, working at the different frequency bands (i.e., 5 GHz and 6 GHz). We consider both UL and DL full buffer traffic type, which means that there is always data to send, so it is often used to evaluate the maximum throughput of the network. The complete set of parameters used in the simulation is shown in Table I.
We first analyzed the simulation results of adding one or two EHT R1 features and compare them with the results of IEEE 802.11ax (i.e., 256 MPDU aggregation for A-MPDU, one link with 160 MHz bandwidth, 1K-QAM). Although the simulation results with one or two R1 features are better than the 802.11ax performance, no single individual feature can reach 30 Gbps, and only combining 320 MHz bandwidth with 2 MLO links for UL MU can achieve this. It should be noted that this part of results figures are omitted due to a lack of space.
Second, we consider the combination of three or four EHT R1 features. The settings of simulation parameters are described in Table II and the activated R1 features are marked with checkmarks. It is noteworthy that there are two different ways to combine 320 MHz bandwidth links. The simulation results are shown in Fig. 3 and the throughput exceeding 30 Gbps is bolded. It can be seen that with three combined R1 features, case 1-3 for UL MU, case 1-6 for MU cases achieve 30 Gbps; with four combined R1 features, case 2-1 for MU cases and case 2-2 for both MU and SU cases achieves 30 Gbps. In addition, it can be concluded that multiple links with 320 MHz can improve throughput. For example, case 2-2 with four R1 features has 262.77%, 270.36%, 373.70%, 401.76% throughput gain in DL SU, UL SU, DL SU and UL MU respectively compared with 802.11ax. In case 3, three links are applied by adding one link with 160 MHz based on case 2-1 and it achieves 30 Gbps for both MU and SU cases. Therefore, adding links can improve the throughput.
@@ V Discussion
In this section, we briefly summarize and discuss the simulation results of low latency and throughput, which is the key objectives of 802.11be. Our simulations show how much throughput can be achieved by combining R1 features. Combining all four R1 features achieves 30 Gbps for both SU and MU cases. Multi-link with two links using 320 MHz channels and 1024 aggregation for the UL MU case can also achieve 30 Gbps throughput. Generally, the impact on throughput for EHT features from lowest to highest is: 4K-QAM, 1024 aggregation, 320 MHz, multi-link. In addition, our simulations show how much delay improvement can be achieved by applying MLO. We compare 802.11ax with 2 links and 802.11be with 2 links, and each link has the same bandwidth. In this case, the delay of 802.11be is reduced around 30% compared to the delay of 802.11ax. In addition, compared with SU-MIMO, OFDMA can significantly reduce latency because of its bandwidth resource allocation scheme and flexible scheduling strategy.
@@ VI Conclusion
Wi-Fi 7 will lead to extremely high throughput and low latency communications, revolutionizing real-time applications through digital augmentation. In this article, we detail the recently updated timeline and key technologies that IEEE 802.11be has adopted on Wi-Fi 7 PHY and MAC, which dominate the improvement of Wi-Fi 7 performance. We demonstrate that the R1 features of 802.11be can support 30 Gbps throughput and low latency through standard-compliant simulations. Currently, the work on Wi-Fi 7 is almost complete, and the final amendment is scheduled to be released in December 2024. Looking ahead, further research will continue to be studied to lay a foundation for the development of the next generation of Wi-Fi.

### v3

@@ Abstract
As emerging applications demand increasingly higher throughput, IEEE standard 802.11be – Extremely High Throughput (EHT), also known as Wi-Fi 7, was published on July 22, 2025. It can be used to meet the demand for the throughput of 4K/8K videos up to tens of Gbps and low-latency video applications such as virtual reality (VR) and augmented reality (AR). Wi-Fi 7 not only scales Wi-Fi 6 with doubled bandwidth, but also supports real-time applications, which brings revolutionary changes to Wi-Fi. In this article, we start by introducing the main objectives and timeline of Wi-Fi 7 and then list the latest key techniques which promote the performance improvement of Wi-Fi 7. Finally, we validate the most critical objectives of Wi-Fi 7 – the potential up to 30 Gbps throughput and lower latency. System-level simulation results suggest that by combining the new techniques, Wi-Fi 7 achieves 30 Gbps throughput and lower latency than Wi-Fi 6.
@@ I Introduction
In this article, we introduce the application scenarios of the next-generation Wi-Fi and the main objectives of IEEE 802.11be, and summarize the recently launched standardization timeline. We delved into the latest concrete features to be adopted in the 802.11be amendment and detailed the implementations and benefits of each feature. In addition, we discuss two important and appealing issues – latency and throughput. In particular, we provide system-level simulation results to show that the features defined by 802.11be can achieve the throughput of 30 Gbps and improve latency performance. Altogether, this article aims to summarize Wi-Fi 7 and provide a guide for researchers and general audiences.
@@ II Wi-Fi 7 Progress
The standardization and certification timeline of Wi-Fi 7 is illustrated in Fig. 1. The IEEE 802.11 WG approved the formation of a Topic Interest Group (TIG) in May 2018, and further established a Study Group (SG) and a Task Group (TG) in July 2018 and May 2019 respectively. The primary objective of these groups is to ensure the rapid and stable progress of 802.11be EHT.
In September 2020, 802.11be draft 0.1 was released, followed by draft 1.0 and draft 2.0 in May 2021 and May 2022, respectively. The IEEE 802.11be D7.0 draft, which is the final amendment, was approved by the IEEE 802 Executive Committee (802 EC) in September 2024, and the draft has now been frozen, marking the completion of the standard’s development phase. Since early 2024, the Wi-Fi Alliance has also begun the certification process for Wi-Fi 7. In July 2025, IEEE standard 802.11be was officially published. The 802.11be amendment spans five years and is expected to provide significant performance enhancements for Wi-Fi 7 users in all respects.
@@
Figure 1: Illustration of recent Wi-Fi standardization and certification timeline.
@@ III Wi-Fi 7 Features
Spectrum is the cornerstone of wireless communication, and the emergence of new generation of wireless communication technologies usually accompanies the use of new spectrum bands. However, the obtainable bandwidth over 2.4 GHz and 5 GHz frequency bands cannot meet the throughput and latency requirements of emerging service such as AR/VR and online gaming [2]. EHT expands its bandwidth by aggregating spectrum across the 2.4 GHz, 5 GHz, and 6 GHz bands, achieving a total channel width of up to 320 MHz, which doubles the maximum nominal throughput compared to 802.11ax.
@@ III-E Multi-Link Operations (MLO)
The MLO is an important new feature of 802.11be [10]. Through MLO, APs and STAs will be provided with the capabilities to transmit and receive data from the same traffic flow over multiple radio interfaces. Its adoption represents the development tendency towards multi-link communications. TGbe has made changes in architecture, management functions, and power save to implement MLO and utilizes all available spectrum resources in 2.4 GHz, 5 GHz and 6 GHz. Next, we explore the key features of MLO.
@@ III-E1 Architecture
802.11be defines the multi-link device (MLD), which incorporates the traditional single-link AP or STA functionalities as internal affiliated entities across multiple links. Both AP MLDs and STA MLDs are single devices with multiple physical radio interfaces. However, it is noteworthy that there is only one MAC address per MLD, and individual parameters for each interface are not lost. Fig. 2 shows the MLD architecture, multiple connections can be set up between APs and STAs. MLD can be considered as a collaborative work scheme. This scheme allows retransmission of packets on any link, which simplifies dynamic link switching and packet reassembly.
MLD is divided into two different transmission modes: simultaneous TX and RX (STR) mode and non-STR (NSTR) mode [11]. STR MLD can transmit frames asynchronously over multiple links. This mode can be applied to both AP MLD and non-AP MLD and keep independent channel access in different links. STR MLD can achieve higher throughput, but power consumption may increase due to simultaneous operation of multiple interfaces. Therefore, this mode must be followed by a power save mechanism, which will be described later. NSTR MLD does not allow information to be transmitted over an idle interface while they are receiving through another and and can only be applied to non-AP MLDs. This mode can avoid in-device coexistence (IDC) interference problems, but at the cost of reducing throughput.
@@ III-E3 Power Save
Since the most major source of traffic in the information age is portable mobile devices, issues related to power consumption must be carefully considered. The use of MLD results in the presence of at least two devices at the same time, which consumes twice more energy. When the traffic is light, the device does not need to continuously listen to more than one link. The basic power management mechanism defines two working modes for Wi-Fi devices: active mode and power save (PS) mode [13]. In active mode, the device normally transmits and receives frames. In PS mode, the device alternates between the awake state and the doze state to conserve power.
The target wake time (TWT) mechanism was first introduced in IEEE 802.11ah, underwent significant enhancements in IEEE 802.11ax, and was further extended in IEEE 802.11be specifically to support MLO scenarios. The key part of TWT is the initial negotiation mechanism, in which the TWT session parameters are determined [12]. Through the TWT, the STA may exchange frames with the AP during a period of waking up (TWT service periods), and keep dozing out of the TWT service periods. Under the MLO architecture, an MLD may negotiate individual TWT agreements with a peer MLD. Note that the individual TWT agreement is negotiated between the STAs affiliated with the MLDs that are operating on an enabled link and is not negotiated between two MLDs.
@@ III-F Low-Latency Operations
TABLE II: Simulation settings with three or four features
@@
Figure 3: Simulation results of 802.11be throughput (with three or four features).
@@ IV-A PAR throughput verification
In the validation process for achieving 30 Gbps throughput, four features were considered to meet the objectives outlined in the PAR for the project. These features included multi-link support (with 2 links and STR transmission mode), 4K-QAM modulation, aggregation of 1024 MPDU aggregation per A-MPDU, and a 320 MHz bandwidth. These features collectively ensure a guaranteed 30 Gbps throughput for both single user (SU) and multi-user (MU) scenarios. In the simulation scenario, each BSS consists of an AP MLD and $N$ non-AP MLDs placed around it and the throughput is measured at MAC data service access point (SAP). In terms of simulation setup parameters, the number of non-AP MLD is 8 and the size of area is 20 m $\times$ 20 m. The TX power of AP MLD is 20 dBm and the TX power of non-AP MLD is 15 dBm. The number of links is 2, working at the different frequency bands (i.e., 5 GHz and 6 GHz). We consider both UL and DL full buffer traffic type, which means that there is always data to send, so it is often used to evaluate the maximum throughput of the network. The complete set of parameters used in the simulation is shown in Table I.
We first analyzed the simulation results of adding one or two EHT features and compare them with the results of IEEE 802.11ax (i.e., 256 MPDU aggregation for A-MPDU, one link with 160 MHz bandwidth, 1K-QAM). Although the simulation results with one or two features are better than the 802.11ax performance, no single individual feature can reach 30 Gbps, and only combining 320 MHz bandwidth with 2 MLO links for UL MU can achieve this. It should be noted that this part of results figures are omitted due to a lack of space.
Second, we consider the combination of three or four EHT features. The settings of simulation parameters are described in Table II and the activated features are marked with checkmarks. It is noteworthy that there are two different ways to combine 320 MHz bandwidth links. The simulation results are shown in Fig. 3 and the throughput exceeding 30 Gbps is bolded. It can be seen that with three combined features, case 1-3 for UL MU, case 1-6 for MU cases achieve 30 Gbps; with four combined features, case 2-1 for MU cases and case 2-2 for both MU and SU cases achieves 30 Gbps. In addition, it can be concluded that multiple links with 320 MHz can improve throughput. For example, case 2-2 with four features has 262.77%, 270.36%, 373.70%, 401.76% throughput gain in DL SU, UL SU, DL SU and UL MU respectively compared with 802.11ax. In case 3, three links are applied by adding one link with 160 MHz based on case 2-1 and it achieves 30 Gbps for both MU and SU cases. Therefore, adding links can improve the throughput.
@@ V Discussion
In this section, we briefly summarize and discuss the simulation results of low latency and throughput, which is the key objectives of 802.11be. Our simulations show how much throughput can be achieved by combining features. Combining all four features achieves 30 Gbps for both SU and MU cases. Multi-link with two links using 320 MHz channels and 1024 aggregation for the UL MU case can also achieve 30 Gbps throughput. Generally, the impact on throughput for EHT features from lowest to highest is: 4K-QAM, 1024 aggregation, 320 MHz, multi-link. In addition, our simulations show how much delay improvement can be achieved by applying MLO. We compare 802.11ax with 2 links and 802.11be with 2 links, and each link has the same bandwidth. In this case, the delay of 802.11be is reduced around 30% compared to the delay of 802.11ax. In addition, compared with SU-MIMO, OFDMA can significantly reduce latency because of its bandwidth resource allocation scheme and flexible scheduling strategy.
@@ VI Conclusion
Wi-Fi 7 will lead to extremely high throughput and low latency communications, significantly enhancing real-time applications through digital augmentation. In this article, we detail the recently updated timeline and key technologies that IEEE 802.11be has adopted on Wi-Fi 7 PHY and MAC, which dominate the improvement of Wi-Fi 7 performance. We demonstrate that the features of 802.11be can support 30 Gbps throughput and low latency through standard-compliant simulations. Looking ahead, further research will continue to be studied to lay a foundation for the development of the next generation of Wi-Fi.